Create a table for the equation:

y= 2x

1
Expert's answer
2021-05-10T18:29:35-0400

Given,

"y=2^x"

Now,

let "x=0,1,2,3,4....."

"x=0"

"y=2^0=1"


"x=1"

"y=2^1= 2"


"x=2"

"y=2^2=4"


"x=3"

"y=2^3 = 8"


"x=4"

"y=2^4 = 16"


"y=1,2, 4, 8,16...."